Problem 1.10 You deposit $5,000 in an account that earns 5% interest compounded annually in years 1 and 2, and thereafter a continuous rate δ(t) 2/(t + 1) (t > 0). What is the value of the account after 5 years?please use make sure you use delta (force of interest) in this equation. It's something like e^integral from year2 to year5 (2/t+1)
Answer #1

Answer:

1St Two Years Interest Rate 5%
And next 3 Years Interest Rate 5.127%(continuous compounding)
Effective annual interest rate in case of continuous compounding is:
EAR = e^rs -1
rate of interest(rs) = 5%
EAR=5.127%
Where e = 2.718281828
Year Op. Deposit Interest rate Interest Closing balance of Deposit
(A) (B) ( C ) D=B*C   E= B+D
1 5000 5%      250.00              5,250.00
2 5250 5%      262.50              5,512.50
3 5,512.50 5.127%      282.63              5,795.13
4 5,795.13 5.127%      297.12              6,092.25
5 6,092.26 5.127%      312.36              6,404.61